# Disentangling Latent Shifts of In-Context Learning Through Self-Training

### v1

@@ 3 Experiments
We utilize a set of decoder-only autoregressive LLMs in our experiments. Specifically, we employ Hugging Face implementations (Wolf et al., 2020) of Llama 3 (8B) (Dubey et al., 2024) and Phi 3 (mini 4k) (Abdin et al., 2024) as our primary models, with additional comparison results for Llama 2 (7B) (Touvron et al., 2023). Detailed information about the models is provided in Table 9 in the Appendix.
Evaluation.
We evaluate the models on the following benchmarks:
  • GLUE (Wang et al., 2018): A standard benchmark for evaluating natural language understanding. We select the following datasets: four binary classification tasks for single sequences (cola, sst, rte), three binary classification tasks for sequence pairs (mrpc, qqp, qnli), and one multi-class classification task for sequence pairs (mnli). We follow the standard practice of evaluating models on the development sets. When evaluating generalization performance, we follow the standard practice and use Matthew’s correlation for cola, $F_{1}$ for mrpc and qqp, and accuracy for the remaining datasets;
  • MMLU (Hendrycks et al., 2021): We evaluate the accuracy of multiple choice question answering on the MMLU benchmark, selecting two datasets with a sufficient number of instances for robust evaluation: “elementary math” (math), assessing basic mathematical reasoning skills, and “miscellaneous” (misc), which covers diverse topics.
In our evaluation, we compute the first-token probability of the task verbalizers. We design the prompt template to guide the model toward generating the answer within the first token and limit the predictions to a subset of verbalizers (cf. Appendix E for details on prompt templates).
Baselines and Methods.
We evaluate sticl by comparing it against three baselines and two ICL disentanglement methods:
  • Zero-Shot ($\mathbf{0}$-shot): Predictions made without any demonstrations;
  • Standard ICL ($\mathbf{n}$-shot): Utilizes $n$ demonstrations as context during inference;
  • Pattern-Based Fine-Tuning (PBFT) (Schick & Schütze, 2021): Fine-tunes the model using patterns learned from data, framed as a language modeling task. In our experiments, we fine-tune an adapter module instead of the whole LLM;
  • In-Context Vectors (ICV) (Liu et al., 2023): A forward pass is used on demonstration examples to create in-context vectors from the hidden states of the LLM;
  • Batch-ICL (Zhang et al., 2024): Utilizes multiple separate one-shot forward computations and aggregates the resulting meta-gradients based on the attention head outputs.
In the experiments, we use $n\in\{4,8,16,32\}$ instances for demonstrations and compare methods using a fixed number of demonstrations. Unless stated otherwise, we run each experiment $10$ times with different seeds, which select different demonstrations in each run. In addition to the generalization scores, we report the standard deviation of the runs as an indicator of method stability. We evaluate performance on the GLUE development sets, while for the MMLU datasets, we sample $200$ instances for evaluation.
sticl variants.
We employ three variants of sticl, which differ in the variability of demonstrations they use, either in terms of selection or ordering:
  • sticl-Fixed (sticl-F): Uses a fixed set of demonstrations throughout training;
  • sticl-Shuffle (sticl-S): Shuffles the order of demonstrations at the start of each epoch;
  • sticl-Resample (sticl-R): Randomly resamples demonstrations before each epoch. [^4] Although sticl-R uses the same number of demonstrations during inference as the other approaches, it requires access to a larger pool of labeled data since it draws new demonstrations in each epoch.
@@ 3.1 Generalization and Stability
GLUE MMLU Model Method rte sst qnli mnli cola mrpc qqp math misc Llama 3 (8B) 0-shot $62.3$ $79.1$ $64.3$ $59.9$ $44.6$ $63.6$ $61.1$ $31.5$ $62.5$ $n$-shot $75.1_{6.5}$ $93.5_{2.0}$ $77.0_{5.5}$ $68.0_{3.0}$ $58.5_{4.0}$ $74.0_{2.5}$ $70.0_{3.0}$ $43.5_{3.5}$ $84.0_{4.0}$ PBFT $73.2_{3.8}$ $93.8_{1.5}$ $77.8_{6.0}$ $67.4_{3.5}$ $56.5_{3.0}$ $72.0_{2.0}$ $68.0_{2.5}$ $44.0_{3.8}$ $83.5_{4.5}$ ICV $72.9_{2.7}$ $92.2_{1.8}$ $74.5_{6.3}$ $67.0_{4.2}$ $57.3_{3.5}$ $73.4_{2.3}$ $69.1_{2.8}$ $41.5_{4.3}$ $67.0_{4.2}$ Batch-ICL $77.8_{4.7}$ $94.1_{2.2}$ $78.0_{6.0}$ $70.9_{3.5}$ $59.8_{3.7}$ $75.2_{2.2}$ $\underline{72.5}_{2.7}$ $36.2_{4.0}$ $81.0_{2.5}$ sticl-F $83.4_{0.3}$ $95.1_{\textbf{0.6}}$ $\underline{80.3}_{\textbf{1.4}}$ $72.1_{2.5}$ $\underline{63.7}_{\textbf{1.5}}$ $76.2_{1.8}$ $71.9_{1.9}$ $\underline{46.0}_{2.3}$ $\underline{86.0}_{2.3}$ sticl-S $\underline{86.0}_{\textbf{0.6}}$ $\textbf{96.1}_{1.2}$ $\textbf{81.4}_{2.2}$ $\underline{73.1}_{\textbf{2.0}}$ $\textbf{64.3}_{2.2}$ $\textbf{77.7}_{\textbf{1.5}}$ $\textbf{73.1}_{\textbf{1.8}}$ $\textbf{49.5}_{\textbf{2.0}}$ $\textbf{88.0}_{\textbf{2.2}}$ sticl-R $\textbf{86.5}_{3.0}$ $\underline{95.5}_{0.8}$ $79.0_{4.3}$ $\textbf{73.5}_{3.0}$ $62.5_{2.8}$ $\underline{76.5}_{1.9}$ $72.0_{2.2}$ $44.0_{2.7}$ $85.5_{3.3}$ Phi 3 (mini 4k) 0-shot $60.6$ $78.3$ $61.1$ $58.1$ $43.7$ $63.1$ $57.8$ $29.5$ $52.0$ $n$-shot $72.1_{5.2}$ $90.6_{2.1}$ $75.6_{3.2}$ $65.3_{3.1}$ $55.5_{4.1}$ $71.1_{2.6}$ $66.2_{3.7}$ $37.5_{3.6}$ $75.5_{4.1}$ PBFT $70.6_{4.3}$ $90.9_{1.9}$ $73.6_{3.4}$ $63.6_{3.6}$ $53.6_{3.1}$ $69.6_{2.3}$ $64.6_{2.6}$ $36.5_{4.1}$ $73.5_{4.6}$ ICV $71.5_{3.1}$ $89.1_{2.1}$ $74.3_{3.2}$ $64.1_{4.1}$ $54.1_{3.6}$ $70.8_{2.4}$ $65.4_{2.9}$ $36.0_{4.6}$ $74.0_{4.3}$ Batch-ICL $75.3_{4.2}$ $91.2_{2.6}$ $76.6_{3.1}$ $67.1_{3.6}$ $56.1_{4.1}$ $72.6_{2.6}$ $67.3_{2.8}$ $38.0_{3.9}$ $76.0_{4.1}$ sticl-F $\underline{80.4}_{\textbf{1.2}}$ $92.1_{\textbf{1.6}}$ $\underline{78.2}_{\textbf{1.3}}$ $\underline{69.7}_{2.4}$ $\underline{59.5}_{2.5}$ $73.5_{2.1}$ $\underline{68.6}_{2.2}$ $\underline{40.5}_{3.2}$ $\underline{77.5}_{3.6}$ sticl-S $\textbf{82.4}_{1.1}$ $\textbf{93.2}_{\textbf{1.6}}$ $79.2_{1.4}$ $\textbf{70.4}_{\textbf{1.1}}$ $\textbf{60.7}_{\textbf{2.3}}$ $\textbf{74.1}_{\textbf{1.4}}$ $\textbf{69.6}_{\textbf{1.9}}$ $\textbf{41.5}_{\textbf{2.3}}$ $\textbf{78.0}_{\textbf{3.3}}$ sticl-R $79.0_{1.9}$ $\underline{92.6}_{2.0}$ $\textbf{79.6}_{2.9}$ $68.6_{3.9}$ $58.6_{2.9}$ $\underline{73.6}_{2.0}$ $68.1_{2.3}$ $39.5_{3.6}$ $77.0_{3.7}$
@@
We first evaluate the generalization and stability of STICL on ID data. Table 1 reports the 16-shot ID generalization scores along with standard deviations. Across all datasets and models, sticl-S consistently achieves the best generalization scores, outperforming standard ICL, PBFT, and the disentanglement methods ICV and Batch-ICL (cf. Table 5 in the Appendix for results with Llama 2). Compared to standard ICL, sticl-S shows absolute improvements ranging from $2.6\%$ to $11.9\%$ for Llama 3 and $2.5\%$ to $10.3\%$ for Phi 3, where the differences in scores are statically significant across all datasets. [^5] We assess the statistical significance using a two-tailed Wilcoxon signed-rank test ($p<0.05$), applying the Holm-Bonferroni method for family-wise error rate correction due to multiple comparisons. Similar patterns hold for $n\in\{4,8,32\}$, where sticl-S also surpasses standard ICL (cf. Table 6 in the Appendix for other $n$-shot setups). Additionally, when a larger set $\mathcal{D}_{\text{unlab}}$ is used, there is a marginal improvement in scores, while stability improves even further (cf. Table 7 in the Appendix). Notably, the improvements in generalization with sticl-S, compared to standard ICL – the teacher model in sticl– provide strong evidence that the student model is exhibiting weak-to-strong generalization; we provide a more detailed analysis of this phenomenon in Section 4. While the sticl-F and sticl-R variants also show similar generalization scores as sticl-S, they generally exhibit higher variance compared to sticl-S, making sticl-S the preferred choice due to its higher stability with respect to demonstration selection – it improves upon standard $n$-shot ICL across all datasets and models. This is supported by the statistically significant differences in standard deviations on all datasets for Llama 3 and on all but qnli for Phi 3.666 We evaluate significance using a two-tailed Levene’s test ($p<0.05$), applying the Holm-Bonferroni method for family-wise error rate correction.
@@
Having looked at stability with respect to demonstration selection, we now turn to a more focused evaluation of stability with respect to demonstration ordering. Table 2 reports the standard deviations across 50 runs, where the same set of demonstrations is used, but their order is shuffled for each run. Designed to adapt to shuffled demonstrations, sticl-S shows the highest stability to demonstration ordering, as evidenced by the smallest standard deviation. The stability improvements with sticl-S over standard ICL are statistically significant across all datasets.6
@@
We next assess the capacity of STICL to perform OOD generalization by fine-tuning an adapter on one dataset and then applying the student model to a different dataset within the same task category, simulating a near-OOD scenario with pairs of closely related datasets. Table 3 shows the OOD generalization scores for such pairs of datasets in the GLUE benchmark. The results show that sticl-S not only outperforms other methods in OOD generalization but also maintains higher stability when adapting to new domains (cf. Table 8 in the Appendix for results with other models).
@@
Table 2: Standard deviations of generalization scores across $50$ runs with different demonstration ordering in the $16$-shot setup. The smallest standard deviations for each dataset are highlighted in bold, while the second-smallest are underlined.
@@
GLUE MMLU Model Method rte sst qnli mnli cola mrpc qqp math misc LLama 3 (8B) $n$-shot $4.81$ $1.62$ $4.19$ $2.22$ $3.04$ $1.81$ $2.03$ $2.52$ $2.87$ PBFT $2.71$ $1.14$ $4.53$ $2.69$ $2.27$ $1.57$ $1.82$ $2.70$ $3.22$ ICV $2.09$ $1.23$ $4.08$ $2.81$ $1.95$ $1.61$ $2.03$ $1.96$ $3.18$ Batch ICL $3.04$ $1.47$ $2.89$ $2.24$ $2.53$ $\underline{1.42}$ $1.74$ $2.51$ $2.59$ sticl-F $\underline{1.32}$ $\underline{0.72}$ $\underline{1.53}$ $\underline{1.83}$ $\underline{1.76}$ $1.54$ $\underline{1.38}$ $\underline{1.89}$ $\underline{2.07}$ sticl-S 0.22 0.53 1.04 1.21 1.28 0.73 1.14 1.22 0.97 sticl-R $2.04$ $1.34$ $2.47$ $2.05$ $1.85$ $1.48$ $1.64$ $2.03$ $2.51$
@@
Table 3: OOD generalization scores for Llama 2 and Phi 3 averaged over 10 runs, with standard deviations shown as subscripts. For each dataset pair, demonstrations are taken from the left dataset, and the model is tested on the right dataset. Columns represent results on the right datasets. The highest scores and lowest standard deviations are in bold, and the second-highest scores are underlined.
Model Method qnli $\rightarrow$ rte rte $\rightarrow$ qnli qqp $\rightarrow$ mrpc mrpc $\rightarrow$ qqp Llama 3 (8B) $n$-shot $66.3_{2.4}$ $69.6_{1.3}$ $66.5_{1.9}$ $62.2_{2.3}$ PBFT $66.1_{1.5}$ $69.1_{1.6}$ $67.2_{1.8}$ $62.4_{1.2}$ ICV $65.7_{1.2}$ $68.7_{2.3}$ $67.5_{1.6}$ $63.0_{2.1}$ Batch-ICL $65.3_{1.4}$ $66.3_{2.5}$ $64.9_{2.3}$ $62.1_{2.1}$ sticl-F $\underline{67.5}_{1.1}$ $\underline{70.5}_{1.4}$ ${68.5}_{\textbf{1.0}}$ ${64.4}_{1.5}$ sticl-S $\textbf{69.0}_{\textbf{0.5}}$ $\textbf{71.3}_{\textbf{0.7}}$ $\textbf{69.0}_{2.2}$ $\underline{66.4}_{\textbf{1.1}}$ sticl-R $67.1_{1.7}$ $70.0_{1.4}$ $68.0_{2.7}$ $\textbf{68.3}_{2.0}$

### v2

@@ 3.1 Experimental Setup
We perform our experiments using decoder-only autoregressive language models provided by Hugging Face [40]. Specifically, we employ Llama 3 (8B) [11] and Phi 3 (mini 4k) [1], along with Llama 2 (7B) [34] for comparative purposes. Further details about the models are listed in Table 11 of the Appendix.
We assess model performance on seven tasks from the GLUE benchmark [36], covering single-sequence binary classification (cola, sst, rte), sequence-pair binary classification (mrpc, qqp, qnli), and sequence-pair multi-class classification (mnli). Evaluation metrics follow established standards: Matthew’s correlation for cola, $F_{1}$ scores for mrpc and qqp, and accuracy for the remaining tasks, with evaluations conducted on the development sets. Additionally, we measure accuracy on selected datasets from the MMLU benchmark [13], specifically “elementary math” (math) and “miscellaneous” (misc). Predictions are made based on the probability of generating specific verbalizer tokens as the first token output by the models, facilitated by carefully crafted prompts designed explicitly for single-token answers (see Appendix F for detailed templates).
Our experiments compare wilda with several baselines, including Zero-Shot ($\mathbf{0}$-shot) inference, which generates predictions without demonstrations, Standard ICL ($\mathbf{n}$-shot), which uses $n$ demonstrations at inference time, and Pattern-Based Fine-Tuning (PBFT) [32], which fine-tunes an adapter module on data-specific patterns. We also include two ICL disentanglement methods, In-Context Vectors (ICV) [27], which leverages hidden-state representations from demonstration examples, and Batch-ICL [41], which aggregates meta-gradients across multiple one-shot runs. All methods are evaluated using a fixed number of demonstrations, with $n\in\{4,8,16,32\}$. Each experiment is repeated 10 times with different random seeds, resulting in varied demonstration selections across runs. Alongside generalization scores, we report the standard deviation across these runs as an indicator of each method’s stability. Evaluations for GLUE are conducted on the development sets, whereas for the MMLU datasets, we randomly sample 200 instances for evaluation.
We employ three variants of wilda, each differing in how demonstrations are selected or ordered during training: Fixed (wilda-F) uses a fixed, unchanging set of demonstrations throughout the entire training process, Shuffle (wilda-S) uses the same demonstrations throughout training but shuffles their order at the beginning of each epoch, and Resample (wilda-R) draws a new set of demonstrations from a larger labeled pool at each epoch.
@@ 3.2 Generalization and Stability
GLUE MMLU Model Method rte sst qnli mnli cola mrpc qqp math misc Llama 3 (8B) 0-shot $62.3$ $79.1$ $64.3$ $59.9$ $44.6$ $63.6$ $61.1$ $31.5$ $62.5$ $n$-shot $75.1_{6.5}$ $93.5_{2.0}$ $77.0_{5.5}$ $68.0_{3.0}$ $58.5_{4.0}$ $74.0_{2.5}$ $70.0_{3.0}$ $43.5_{3.5}$ $84.0_{4.0}$ PBFT $73.2_{3.8}$ $93.8_{1.5}$ $77.8_{6.0}$ $67.4_{3.5}$ $56.5_{3.0}$ $72.0_{2.0}$ $68.0_{2.5}$ $44.0_{3.8}$ $83.5_{4.5}$ ICV $72.9_{2.7}$ $92.2_{1.8}$ $74.5_{6.3}$ $67.0_{4.2}$ $57.3_{3.5}$ $73.4_{2.3}$ $69.1_{2.8}$ $41.5_{4.3}$ $67.0_{4.2}$ Batch-ICL $77.8_{4.7}$ $94.1_{2.2}$ $78.0_{6.0}$ $70.9_{3.5}$ $59.8_{3.7}$ $75.2_{2.2}$ $\underline{72.5}_{2.7}$ $36.2_{4.0}$ $81.0_{2.5}$ wilda-F $83.4_{0.3}$ $95.1_{\textbf{0.6}}$ $\underline{80.3}_{\textbf{1.4}}$ $72.1_{2.5}$ $\underline{63.7}_{\textbf{1.5}}$ $76.2_{1.8}$ $71.9_{1.9}$ $\underline{46.0}_{2.3}$ $\underline{86.0}_{2.3}$ wilda-S $\underline{86.0}_{\textbf{0.6}}$ $\textbf{96.1}_{1.2}$ $\textbf{81.4}_{2.2}$ $\underline{73.1}_{\textbf{2.0}}$ $\textbf{64.3}_{2.2}$ $\textbf{77.7}_{\textbf{1.5}}$ $\textbf{73.1}_{\textbf{1.8}}$ $\textbf{49.5}_{\textbf{2.0}}$ $\textbf{88.0}_{\textbf{2.2}}$ wilda-R $\textbf{86.5}_{3.0}$ $\underline{95.5}_{0.8}$ $79.0_{4.3}$ $\textbf{73.5}_{3.0}$ $62.5_{2.8}$ $\underline{76.5}_{1.9}$ $72.0_{2.2}$ $44.0_{2.7}$ $85.5_{3.3}$ Phi 3 (mini 4k) 0-shot $60.6$ $78.3$ $61.1$ $58.1$ $43.7$ $63.1$ $57.8$ $29.5$ $52.0$ $n$-shot $72.1_{5.2}$ $90.6_{2.1}$ $75.6_{3.2}$ $65.3_{3.1}$ $55.5_{4.1}$ $71.1_{2.6}$ $66.2_{3.7}$ $37.5_{3.6}$ $75.5_{4.1}$ PBFT $70.6_{4.3}$ $90.9_{1.9}$ $73.6_{3.4}$ $63.6_{3.6}$ $53.6_{3.1}$ $69.6_{2.3}$ $64.6_{2.6}$ $36.5_{4.1}$ $73.5_{4.6}$ ICV $71.5_{3.1}$ $89.1_{2.1}$ $74.3_{3.2}$ $64.1_{4.1}$ $54.1_{3.6}$ $70.8_{2.4}$ $65.4_{2.9}$ $36.0_{4.6}$ $74.0_{4.3}$ Batch-ICL $75.3_{4.2}$ $91.2_{2.6}$ $76.6_{3.1}$ $67.1_{3.6}$ $56.1_{4.1}$ $72.6_{2.6}$ $67.3_{2.8}$ $38.0_{3.9}$ $76.0_{4.1}$ wilda-F $\underline{80.4}_{\textbf{1.2}}$ $92.1_{\textbf{1.6}}$ $\underline{78.2}_{\textbf{1.3}}$ $\underline{69.7}_{2.4}$ $\underline{59.5}_{2.5}$ $73.5_{2.1}$ $\underline{68.6}_{2.2}$ $\underline{40.5}_{3.2}$ $\underline{77.5}_{3.6}$ wilda-S $\textbf{82.4}_{1.1}$ $\textbf{93.2}_{\textbf{1.6}}$ $79.2_{1.4}$ $\textbf{70.4}_{\textbf{1.1}}$ $\textbf{60.7}_{\textbf{2.3}}$ $\textbf{74.1}_{\textbf{1.4}}$ $\textbf{69.6}_{\textbf{1.9}}$ $\textbf{41.5}_{\textbf{2.3}}$ $\textbf{78.0}_{\textbf{3.3}}$ wilda-R $79.0_{1.9}$ $\underline{92.6}_{2.0}$ $\textbf{79.6}_{2.9}$ $68.6_{3.9}$ $58.6_{2.9}$ $\underline{73.6}_{2.0}$ $68.1_{2.3}$ $39.5_{3.6}$ $77.0_{3.7}$
@@
We first evaluate the generalization and stability of wilda on ID data. Table 1 reports the 16-shot ID generalization scores along with standard deviations. Across all datasets and models, wilda-S consistently achieves the best generalization scores, outperforming standard ICL, PBFT, and the disentanglement methods ICV and Batch-ICL (cf. Table 5 in the Appendix for results with Llama 2). Compared to standard ICL, wilda-S shows absolute improvements ranging from $2.6\%$ to $11.9\%$ for Llama 3 and $2.5\%$ to $10.3\%$ for Phi 3, where the differences in scores are statistically significant across all datasets. [^2] We assess the statistical significance using a two-tailed Wilcoxon signed-rank test ($p<0.05$), applying the Holm-Bonferroni method for family-wise error rate correction. Similar patterns hold for $n\in\{4,8,32\}$, where wilda-S also surpasses standard ICL (cf. Table 6 in the Appendix for other $n$-shot setups). Additionally, when a larger set $\mathcal{D}_{\text{unlab}}$ is used, there is a marginal improvement in scores, while stability improves even further (see Table 7 in the Appendix). Notably, the improvements in generalization with wilda-S, compared to standard ICL (the teacher model in wilda), provide strong evidence that the student model is exhibiting W2S generalization; we provide a more detailed analysis of this phenomenon in Section 4. While the wilda-F and wilda-R variants show similar generalization scores to wilda-S, they typically exhibit higher variance. This makes wilda-S the preferred choice due to its greater stability with respect to demonstration selection, as it consistently improves upon standard $n$-shot ICL across all datasets and models. This is supported by the statistically significant differences in standard deviations on all datasets for Llama 3 and on all but qnli for Phi 3.333We test for significance using a two-tailed Levene’s test ($p<0.05$) and apply the Holm-Bonferroni method to correct for family-wise error rate.
@@
Having looked at stability with respect to demonstration selection, we now turn to a more focused evaluation of stability with respect to demonstration ordering. Table 2 reports the standard deviations across 50 runs, where the same set of demonstrations is used, but their order is shuffled for each run. Designed to adapt to shuffled demonstrations, wilda-S shows the highest stability to demonstration ordering, as evidenced by the smallest standard deviation. The stability improvements with wilda-S over standard ICL are statistically significant across all datasets.3
@@
We next assess the capacity of wilda to perform OOD generalization by fine-tuning an adapter on one dataset and then applying the student model to a different dataset within the same task category, simulating a near-OOD scenario with pairs of closely related datasets. Table 3 shows the OOD generalization scores for such pairs of datasets in the GLUE benchmark. The results show that wilda-S not only outperforms other methods in OOD generalization but also maintains higher stability when adapting to new domains (cf. Table 8 in the Appendix for results with other models).
@@
Table 2: Standard deviations of generalization scores across $50$ runs with varied orderings of $16$ demonstrations. The smallest deviations are in bold, and the second-smallest are underlined.
@@
GLUE MMLU Model Method rte sst qnli mnli cola mrpc qqp math misc LLama 3 (8B) $n$-shot $4.81$ $1.62$ $4.19$ $2.22$ $3.04$ $1.81$ $2.03$ $2.52$ $2.87$ PBFT $2.71$ $1.14$ $4.53$ $2.69$ $2.27$ $1.57$ $1.82$ $2.70$ $3.22$ ICV $2.09$ $1.23$ $4.08$ $2.81$ $1.95$ $1.61$ $2.03$ $1.96$ $3.18$ Batch ICL $3.04$ $1.47$ $2.89$ $2.24$ $2.53$ $\underline{1.42}$ $1.74$ $2.51$ $2.59$ wilda-F $\underline{1.32}$ $\underline{0.72}$ $\underline{1.53}$ $\underline{1.83}$ $\underline{1.76}$ $1.54$ $\underline{1.38}$ $\underline{1.89}$ $\underline{2.07}$ wilda-S 0.22 0.53 1.04 1.21 1.28 0.73 1.14 1.22 0.97 wilda-R $2.04$ $1.34$ $2.47$ $2.05$ $1.85$ $1.48$ $1.64$ $2.03$ $2.51$
@@
Table 3: OOD generalization scores with $16$ shots averaged over 10 runs, with standard deviations shown as subscripts. For each dataset pair, demonstrations are taken from the left dataset, and the model is tested on the right dataset. Columns represent results on the right datasets. The highest scores and lowest standard deviations are in bold, and the second-highest scores are underlined. Values in parentheses indicate differences from ID performance for the corresponding target dataset.
[…]